OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (6, 2) and has a slope of 3? y =___x +___ (Note: I don't have to simplify I just need to put it in this equation)

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

hey @bruhhh plz help

OpenStudy (anonymous):

Hey! :) So if you're given a point and a slope, you're gonna have to use point-slope form, which is y-y1=m(x-x1) Plug in the coordinates & slope like this: y-2=3(x-6) Do you know what to do from there? :)

OpenStudy (anonymous):

i don't have to simplify it

OpenStudy (anonymous):

that's what i don't get

OpenStudy (anonymous):

Wait what, but that's the only way to get to the y=mx+b form?

OpenStudy (anonymous):

ikr

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

this is why i hate math now

OpenStudy (anonymous):

Lol but look Like distributing the slope will make it y-2=3x-18 And move the 2 over: y=3x-16 <--that's your answer

OpenStudy (anonymous):

oh!!

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Haha okay, no prob! :)